# The Covariant Stone-von Neumann Theorem for Locally Compact Quantum Groups

### v1

@@ Abstract.
The Stone-von Neumann Theorem is a fundamental result which unified the competing quantum-mechanical models of matrix mechanics and wave mechanics. It’s mechanism of proof ultimately involved the study of unitary group representations on a Hilbert space. In this article, we continue the broad generalization set out in [HI, modular], analyzing representations of locally compact quantum dynamical systems defined on Hilbert modules, of which the classical result is a special case. We introduce a pair of modular representations which subsume numerous models which appear in the literature, and for certain coactions $(\mathbb{G},A,\alpha)$ recover the multiplicity results of [modular]. As a corollary, we develop a new criterion for identifying strongly regular locally compact quantum groups, related to the study of their dynamics on elementary $C^{*}$-algebras.
@@ 1. Introduction
The fields of science have long benefited from interdisciplinary interactions, and the application of mathematics to these fields is no exception. The past century has arguably witnessed among the most influential discoveries in human history the discovery of quantum phenomena. With its foundation in mathematics and home in physics, the exploration of quantum mechanics has beyond any doubt cast us into the atomic era. Chief in the development of this knowledge were the independently developed models of quantum mechanics, namely Werner Heisenberg’s matrix mechanics and Erwin Schrödinger’s wave mechanics. Where matrix mechanics proposed a governance of subatomic phenomena by certain interacting linear transformations, wave mechanics distinguished itself by the application of partial differential equations. Remarkably, each model yielded the same experimental predictions, suggesting a unity of theory despite cosmetic differences. Using sophisticated new tools developed by Marshall Stone (which today serve as important rudiments in the study of operator algebras), John von Neumann successfully proved the formal mathematical equivalence of these two models in what is now called the Stone-von Neumann Theorem. Loosely stated, the theorem proves that any self-adjoint representation of Heisenberg’s linear operators is merely a multiple of the irreducible representation coming from solutions to Schrödinger’s wave equation. The essential strategy involved associating to a coupled pair of linear transformations two families of linear operators indexed by the topological group $\mathbb{R}^{n}$.
@@
Von Neumann proceeded to study rings of operators, initiating the study of operator algebras which reaches as far in the field of mathematics as quantum mechanics does in physics. Indeed, operator algebraists have developed deep interactions with numerous “classical” mathematical fields, mirroring the drive to unify physics under one “standard model”. Throughout this enrichment of the field, a steady interest in Stone and von Neumann’s previous investigations remained, sustaining a perpetual refinement and generalization of ideas. Major influences include George Mackey and Marc Rieffel, with successive innovations from numerous others. Recently, Ismert and the second author provided a significant innovation, classifying representations of $C^{*}$-dynamical systems of abelian groups represented on Hilbert modules. Quigg and the first two authors expanded this classification to include dynamical systems of general (nonabelian) locally compact groups using divergent techniques from the parallel but distinct lenses of duality for $C^{*}$-actions and coactions.
@@
From the lens of nonabelian duality, the divergence of technique is a necessary cost for the expansion beyond actions of an abelian group. However, over the past few decades the appealing new theory of locally compact quantum groups has emerged to unify these parallel developments under a common theory. In this article, we call upon this theory to integrate the work of Quigg and the first two authors, recovering numerous independent results inspired by Stone and von Neumann’s century-old work. Figuring prominently in this article is the emphasis on corepresentations of locally compact quantum groups, which are a valuable tool for quantum groups which appears implicitly in the coaction setting of [modular]. These objects, other conventions for locally compact quantum groups, and their crossed products are recorded in §2. Heisenberg representations of $C^{*}$-quantum group dynamical systems are defined in §3, where we also explore close relationships with prior evolutionary expressions. Together with these relationships, Theorem 3.9 serves to confirm the proposed definition. In §4, we introduce the Schrödinger representation of a quantum dynamical system, and prove our ultimate result, that every Heisenberg representation of a maximal coaction on an elementary $C^{*}$-algebra is a multiple of the Schrödinger representation. We conclude our investigation with an examination of conditions under which one expects a Stone-von Neumann-type result to hold, which includes results offering new sufficient conditions for maximality of a coaction as well as strong regularity for a locally compact quantum group.
@@ Acknowledgments
The work of JK was partially supported by FWO grant 1246624N. as well as EPSRC grants EP/T03064X/1 and EP/T030992/1.
@@ 2. Preliminaries
In this section we record the conventions we will use throughout the article. All homomorphisms are assumed to be ***-preserving. An elementary $C^{*}$-algebra is a $C^{*}$-algebra $A$ which is isomorphic to the $C^{*}$-algebra of compact operators over some Hilbert space $H$.
@@ 2.2. The Millinery of Locally Compact Quantum Groups
Our main reference for locally compact quantum groups are [kv, buss]. Other standard references include [kustermansuniversal, venchilada, KVvn].
A locally compact quantum group $\mathbb{G}$ is a $C^{*}$-algebra $C_{0}(\mathbb{G})$ together with a nondegenerate ***-homomorphism $\Delta\colon C_{0}(\mathbb{G})\to M(C_{0}(\mathbb{G})\otimes C_{0}(\mathbb{G}))$ called comultiplication which satisfies the coassociativity and density conditions
$\begin{split}(\Delta\otimes\text{id})\Delta&=(\text{id}\otimes\Delta)\Delta,\text{ and}\\ C_{0}(\mathbb{G})=\overline{\operatorname{span}}\{(\omega\otimes\text{id})\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}&=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}.\end{split}$
@@
From $\mathrm{W}^{\mathbb{G}}$, one recovers the $C^{*}$-algebra $C_{0}(\mathbb{G})=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and comultiplication $\Delta(a)=\mathrm{W}^{\mathbb{G}*}(1\otimes a)\mathrm{W}^{\mathbb{G}}\;(a\in C_{0}(\mathbb{G}))$. One can define the dual locally compact quantum group $\widehat{\mathbb{G}}$ — its $C^{*}$-algebra may be recognized as $C_{0}(\widehat{\mathbb{G}})=\overline{\operatorname{span}}\{(\omega\otimes\text{id})(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and the comultiplication is given by $\widehat{\Delta}(a)=\mathrm{W}^{\widehat{\mathbb{G}}*}(1\otimes a)\mathrm{W}^{\widehat{\mathbb{G}}}\;(a\in C_{0}(\widehat{\mathbb{G}}))$, where $\mathrm{W}^{\widehat{\mathbb{G}}}=\chi(\mathrm{W}^{\mathbb{G}*})$. In fact, $\mathrm{W}^{\widehat{\mathbb{G}}}$ is the Kac-Takesaki operator of $\widehat{\mathbb{G}}$. The remaining features of $\widehat{\mathbb{G}}$ will be likewise decorated with hats. One can prove that in fact $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$ and
@@
The culmination of the theory is the recovery of generalized Pontryagin duality $\widehat{\widehat{\mathbb{G}}}=\mathbb{G}$.
@@
Out of a quantum group $\mathbb{G}$, one builds yet another $C^{*}$-algebra, $C_{0}^{u}(\mathbb{G})$, which should be thought of as a universal version of $C_{0}(\mathbb{G})$. It comes with its own comultiplication: a non-degenerate ***-homomorphism $\Delta^{u}\colon C_{0}^{u}(\mathbb{G})\rightarrow M(C_{0}^{u}(\mathbb{G})\otimes C_{0}^{u}(\mathbb{G}))$ and the reducing map, which is a ***-epimorphism $\lambda\colon C^{u}_{0}(\mathbb{G})\rightarrow C_{0}(\mathbb{G})$ respecting comultiplications (i.e. $\Delta\circ\lambda=(\lambda\otimes\lambda)\circ\Delta^{u}$). Furthermore, it has a very useful universal property concerning corepresentations of $\widehat{\mathbb{G}}$ (see Section 2.3 and [kustermansuniversal]).
@@
The final relation states that the opposite of the commutant of $\mathbb{G}$ is unitarily equivalent to $\mathbb{G}$ by the unitary $\widehat{J}J$ (see [KVvn, Section 4]). Each of the quantum groups we consider above admits universal companions, and we will decorate the respective reducing maps using the same notation, e.g. $\hat{\lambda}\colon C_{0}^{u}(\hat{\mathbb{G}})\to C_{0}(\hat{\mathbb{G}})$, and similarly with $\lambda^{\textnormal{op}}$ and $\lambda^{c}$.
@@
2.3. Corepresentations
Given a locally compact quantum group $\mathbb{G}$ and a Hilbert $A$-module $X$, a left corepresentation of $\mathbb{G}$ on $X$ is a unitary $u\in\mathcal{L}(C_{0}(\mathbb{G})\otimes X)$ which satisfies $(\Delta\otimes\text{id})(u)=u_{13}u_{23}$. Likewise, a right corepresentation is a unitary $v\in\mathcal{L}(X\otimes C_{0}(\mathbb{G}))$ which satisfies $(\text{id}\otimes\Delta)(v)=v_{12}v_{13}$. The assignment $u\mapsto u^{\textnormal{op}}=\chi(u)^{*}$ determines a bijective correspondence between left corepresentations of $\mathbb{G}$ and right corepresentations of $\mathbb{G}^{\textnormal{op}}$. The Kac-Takesaki operator is a distinguished left corepresentation of $\mathbb{G}$ on $L^{2}(\mathbb{G})$. In fact, $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$, and regarded this way we call $\mathrm{W}^{\mathbb{G}}$ the left regular corepresentation of $\mathbb{G}$ (on $C_{0}(\widehat{\mathbb{G}})$). We also have the right regular corepresentation of $\mathbb{G}$ (which is indeed a right corepresentation corresponding to a particular left corepresentation of $\mathbb{G}^{\textnormal{op}}$), which is a unitary $\mathrm{V}^{\mathbb{G}}\in M(C_{0}(\widehat{\mathbb{G}}^{\textnormal{c}})\otimes C_{0}(\mathbb{G}))$ satisfying $(\text{id}\otimes\Delta)(\mathrm{V}^{\mathbb{G}})=\mathrm{V}^{\mathbb{G}}_{12}\mathrm{V}^{\mathbb{G}}_{13}$. It also implements the comultiplication, namely $\Delta(x)=\mathrm{V}^{\mathbb{G}}(x\otimes 1)\mathrm{V}^{\mathbb{G}*}$ for $x\in C_{0}(\mathbb{G})$.
@@ 2.3. Corepresentations
2.4. Coaction Crossed Products
Let $A$ be a $C^{*}$-algebra and $\mathbb{G}$ a locally compact quantum group. A left coaction of $\mathbb{G}$ on $A$ is a nondegenerate homomorphism $\alpha\colon A\to M(C_{0}(\mathbb{G})\otimes A)$ which satisfies the following two conditions

### v2

# The Covariant Stone–von Neumann Theorem for Locally Compact Quantum Groups

@@ Abstract.
The Stone–von Neumann Theorem is a fundamental result which unified the competing quantum-mechanical models of matrix mechanics and wave mechanics. In this article, we continue the broad generalization set out by Huang and Ismert and by Hall, Huang, and Quigg, analyzing representations of locally compact quantum dynamical systems defined on Hilbert modules, of which the classical result is a special case. We introduce a pair of modular representations which subsume numerous models available in the literature and, using the classical strategy of Rieffel, prove a Stone–von Neumann-type theorem for maximal actions of regular locally compact quantum groups on elementary C*-algebras. In particular, we generalize the Mackey–Stone–von Neumann Theorem to regular locally compact quantum groups whose trivial actions on $\mathbb{C}$ are maximal and recover the multiplicity results of Hall, Huang, and Quigg. With this characterization in hand, we prove our main result showing that if a dynamical system $(\mathbb{G},A,\alpha)$ satisfies the multiplicity assumption of the generalized Stone–von Neumann theorem, and if the coefficient algebra $A$ admits a faithful state, then the spectrum of the iterated crossed product $\widehat{\mathbb{G}}^{\rm op}\ltimes(\mathbb{G}\ltimes A)$ consists of a single point. In the case of a separable coefficient algebra or a regular acting quantum group, we further characterize features of this system, and thus obtain a partial converse to the Stone–von Neumann theorem in the quantum group setting. As a corollary, we show that a regular locally compact quantum group satisfies the generalized Stone–von Neumann theorem if and only if it is strongly regular.
@@ 1. Introduction
The fields of science have long benefited from interdisciplinary interactions, and the application of mathematics to these fields is no exception. With its foundation in mathematics and home in physics, the exploration of quantum mechanics has beyond any doubt cast us into the atomic era, and chief in the development of this knowledge were the independently developed models of quantum mechanics, namely Werner Heisenberg’s matrix mechanics and Erwin Schrödinger’s wave mechanics. Where matrix mechanics proposed a governance of subatomic phenomena by certain interacting linear transformations, wave mechanics distinguished itself by the application of partial differential equations. Remarkably, each model yielded the same experimental predictions, suggesting a unity of theory despite cosmetic differences. Using sophisticated new tools developed by Marshall Stone (which today serve as important rudiments in the study of operator algebras), John von Neumann successfully proved the formal mathematical equivalence of these two models in what is now called the Stone–von Neumann Theorem. Loosely stated, the theorem proves that any self-adjoint representation of Heisenberg’s linear operators is merely a multiple of the irreducible representation coming from solutions to Schrödinger’s wave equation. The essential strategy involved associating to a coupled pair of linear transformations two families of linear operators indexed by the topological group $\mathbb{R}^{n}$.
@@
Von Neumann proceeded to study rings of operators, initiating the study of operator algebras which reaches as far in the field of mathematics as quantum mechanics does in physics. Indeed, operator algebraists have developed deep interactions with numerous “classical” mathematical fields, mirroring the drive to unify physics under one “standard model”. Throughout this enrichment of the field, a steady interest in Stone and von Neumann’s previous investigations remained, sustaining a perpetual refinement and generalization of ideas. Major influences include George Mackey and Marc Rieffel, with successive advancement from numerous others. Recently, Ismert and the second author provided a significant innovation, classifying representations of $C^{*}$-dynamical systems of abelian groups represented on Hilbert modules. Quigg and the first two authors expanded this classification to include dynamical systems of general (nonabelian) locally compact groups using divergent techniques from the parallel but distinct lenses of duality for $C^{*}$-actions and coactions.
@@
From the perspective of nonabelian duality, the divergence of technique is a necessary cost for the expansion beyond actions of an abelian group. However, recent decades have offered the appealing new theory of locally compact quantum groups to unify these parallel developments under a common theory. In this article, we call upon this theory to integrate the work of Quigg and the first two authors, recovering numerous independent results inspired by Stone and von Neumann’s century-old work, and offering a new analysis of spectral phenomena which may be regarded as a foundation for the multiplicity results which appear in the literature. Figuring prominently in this article is the emphasis on representations of locally compact quantum groups, which are a valuable tool for quantum groups which appears implicitly in the coaction setting of [modular]. These objects, other conventions for locally compact quantum groups, and their crossed products are recorded in §2, together with a new result of general interest to the quantum group community: preservation of the full crossed products of universal quantum groups under duality, which implies under a regularity assumption maximality of the trivial actions. Heisenberg representations of $C^{*}$-quantum group dynamical systems are defined in §3, where we also explore close relationships with prior evolutionary expressions. Together with these relationships, Theorem 3.9 serves to confirm the proposed definition. In §4, for a given dynamical system we introduce the Schrödinger representation, and show that it is a prototype for the Heisenberg representation. For dynamical systems comprising maximal actions and elementary coefficient algebras, we deduce that every Heisenberg representation is a multiple of the Schrödinger representation, and with this characterization in hand, we prove our main result: given a dynamical system $(\mathbb{G},A,\alpha)$ where $\mathbb{G}$ is a locally compact quantum group and $A$ admits a faithful state, if every Heisenberg representation is a multiple of the Schrödinger representation, then the spectrum of the $C^{*}$-algebra $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$ is a one point space. This result is new already for the case of $A=\mathbb{C}$. In this setting, we conclude that (1) if $A$ is separable, then $A$ was an elementary $C^{*}$-algebra all along; and (2) if $\mathbb{G}$ is a regular locally compact group, then the action is maximal. As a final corollary, we show that if $\mathbb{G}$ is regular, the generalized Mackey–Stone–von Neumann Theorem holds for $(\mathbb{G},\mathbb{C},\text{triv})$ if and only if $\mathbb{G}$ is strongly regular (equivalently: the trivial action of $\mathbb{G}$ on $\mathbb{C}$ is maximal).
@@ Acknowledgments
The authors are deeply indebted to the referees, whose commentary vastly improved the exposition of this article.
The work of JK was partially supported by FWO grant 1246624N as well as EPSRC grants EP/T03064X/1 and EP/T030992/1.
@@
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analysed in this study.
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
@@ 2. Preliminaries
In this section we record the conventions we will use throughout the article. All homomorphisms are assumed to be *-preserving. An elementary $C^{*}$-algebra is a $C^{*}$-algebra $A$ which is isomorphic to the $C^{*}$-algebra of compact operators over some Hilbert space $H$.
@@ 2.2. The Millinery of Locally Compact Quantum Groups
Our main references for locally compact quantum groups are [kv, buss]. Other standard references include [kustermansuniversal, venchilada, KVvn].
A locally compact quantum group $\mathbb{G}$ consists of a $C^{*}$-algebra $C_{0}(\mathbb{G})$ together with a nondegenerate *-homomorphism $\Delta\colon C_{0}(\mathbb{G})\to M(C_{0}(\mathbb{G})\otimes C_{0}(\mathbb{G}))$, called comultiplication which satisfies the coassociativity $(\Delta\otimes\text{id})\Delta=(\text{id}\otimes\Delta)\Delta$ and density conditions
$\begin{split}C_{0}(\mathbb{G})&=\overline{\operatorname{span}}\{(\omega\otimes\text{id})\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}\\ &=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)\Delta(x)|~{}\omega\in C_{0}(\mathbb{G})^{*},x\in C_{0}(\mathbb{G})\}.\end{split}$
@@
From $\mathrm{W}^{\mathbb{G}}$, one recovers the $C^{*}$-algebra $C_{0}(\mathbb{G})=\overline{\operatorname{span}}\{(\text{id}\otimes\omega)(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and comultiplication $\Delta(a)=\mathrm{W}^{\mathbb{G}*}(1\otimes a)\mathrm{W}^{\mathbb{G}}\;(a\in C_{0}(\mathbb{G}))$. One defines the dual locally compact quantum group $\widehat{\mathbb{G}}$ — its $C^{*}$-algebra may be recognized as $C_{0}(\widehat{\mathbb{G}})=\overline{\operatorname{span}}\{(\omega\otimes\text{id})(\mathrm{W}^{\mathbb{G}})|~{}\omega\in\mathbb{B}(L^{2}(\mathbb{G}))_{*}\}$ and the comultiplication is given by $\widehat{\Delta}(a)=\mathrm{W}^{\widehat{\mathbb{G}}*}(1\otimes a)\mathrm{W}^{\widehat{\mathbb{G}}}\;(a\in C_{0}(\widehat{\mathbb{G}}))$, where $\mathrm{W}^{\widehat{\mathbb{G}}}=\chi(\mathrm{W}^{\mathbb{G}*})$. In fact, $\mathrm{W}^{\widehat{\mathbb{G}}}$ is the Kac-Takesaki operator of $\widehat{\mathbb{G}}$. The remaining features of $\widehat{\mathbb{G}}$ will be likewise decorated with hats. One can prove that in fact $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$ and
@@
The culmination of the theory is the recovery of generalized Pontryagin duality $\widehat{\widehat{\mathbb{G}}}\cong\mathbb{G}$.
@@
Out of a quantum group $\mathbb{G}$, one builds yet another $C^{*}$-algebra, $C_{0}^{u}(\mathbb{G})$, which should be thought of as a universal version of $C_{0}(\mathbb{G})$. It comes with its own comultiplication: a non-degenerate *-homomorphism $\Delta^{u}\colon C_{0}^{u}(\mathbb{G})\rightarrow M(C_{0}^{u}(\mathbb{G})\otimes C_{0}^{u}(\mathbb{G}))$ and the reducing map, which is a *-epimorphism $\lambda\colon C^{u}_{0}(\mathbb{G})\rightarrow C_{0}(\mathbb{G})$ respecting comultiplications (i.e. $\Delta\circ\lambda=(\lambda\otimes\lambda)\circ\Delta^{u}$). Furthermore, it has a very useful universal property concerning representations of $\widehat{\mathbb{G}}$ which we will encounter in section 2.3 (see also [kustermansuniversal]).
@@
The final relation states that the opposite of the commutant of $\mathbb{G}$ is unitarily equivalent to $\mathbb{G}$ by the unitary $\widehat{J}J$ (see [KVvn, Section 4]). Each of the quantum groups we consider above admits universal companions, and we will decorate the respective reducing maps using the same notation, e.g. $\hat{\lambda}\colon C_{0}^{u}(\widehat{\mathbb{G}})\to C_{0}(\widehat{\mathbb{G}})$, and similarly with $\lambda^{\textnormal{op}}$ and $\lambda^{c}$.
@@
2.3. Representations
Given a locally compact quantum group $\mathbb{G}$ and a Hilbert $A$-module $X$, a left representation of $\mathbb{G}$ (or a left corepresentation of $C_{0}(\mathbb{G})$) on $X$ is a unitary $u\in\mathcal{L}(C_{0}(\mathbb{G})\otimes X)$ which satisfies $(\Delta\otimes\text{id})(u)=u_{13}u_{23}$. Likewise, a right representation is a unitary $v\in\mathcal{L}(X\otimes C_{0}(\mathbb{G}))$ which satisfies $(\text{id}\otimes\Delta)(v)=v_{12}v_{13}$. The assignment $u\mapsto u^{\textnormal{op}}=\chi(u)^{*}$ determines a bijective correspondence between left representations of $\mathbb{G}$ and right representations of $\mathbb{G}^{\textnormal{op}}$. The Kac-Takesaki operator is a distinguished left representation of $\mathbb{G}$ on $L^{2}(\mathbb{G})$. In fact, $\mathrm{W}^{\mathbb{G}}\in M(C_{0}(\mathbb{G})\otimes C_{0}(\widehat{\mathbb{G}}))$, and regarded this way we call $\mathrm{W}^{\mathbb{G}}$ the left regular representation of $\mathbb{G}$ (on $C_{0}(\widehat{\mathbb{G}})$). We also have the right regular representation of $\mathbb{G}$ (which is indeed a right representation corresponding to a particular left representation of $\mathbb{G}^{\textnormal{op}}$), which is a unitary $\mathrm{V}^{\mathbb{G}}\in M(C_{0}(\widehat{\mathbb{G}}^{\textnormal{c}})\otimes C_{0}(\mathbb{G}))$ satisfying $(\text{id}\otimes\Delta)(\mathrm{V}^{\mathbb{G}})=\mathrm{V}^{\mathbb{G}}_{12}\mathrm{V}^{\mathbb{G}}_{13}$. It also implements the comultiplication, namely $\Delta(x)=\mathrm{V}^{\mathbb{G}}(x\otimes 1)\mathrm{V}^{\mathbb{G}*}$ for $x\in C_{0}(\mathbb{G})$.
@@ 2.3. Representations
2.4. Crossed Products
Let $A$ be a $C^{*}$-algebra and $\mathbb{G}$ a locally compact quantum group. A left action of $\mathbb{G}$ on $A$ is a nondegenerate homomorphism $\alpha\colon A\to M(C_{0}(\mathbb{G})\otimes A)$ which satisfies the following two conditions